A ball is launched into the sky at 54.4 feet per second from a 268.8 meter tall building. The equation for the ball’s height, h, at time t seconds is h = -3.2t^2 + 54.4t + 268.8. When will the ball strike the ground?

Found 2 solutions by Alan3354, ikleyn:
Answer by Alan3354(69443)

A ball is launched into the sky at 54.4 feet per second from a 268.8 meter tall building. The equation for the ball’s height, h, at time t seconds is h = -3.2t^2 + 54.4t + 268.8. When will the ball strike the ground?
==============
When h = 0.
-3.2t^2 + 54.4t + 268.8 = 0
======================
Your units don't match.
268.8 meters = 882 feet
----
The commonly used function in English units is
h(t) = -16t^2 + 54.4t + 882
Solve for t when h(t) = 0. Ignore the negative solution.
=====================
the -3.2t^2 term might apply to some other planet, but not Earth.
Earth's acceleration due to gravity is 32 ft/sec/sec.

Answer by ikleyn(52779)

.
Your equation is written with a mistake.

The correct writing is
h = -16t^2 + 54.4t + 268.8.

The ball will strike the ground when h = 0, i.e.
-16t^2 + 54.4t + 268.8 = 0.